OpenStudy (anonymous):

the half-life of a radioactive element is 133 days, but your sample will not be useful to you after 90% of the radioactive nuclei originally present have disintegrated. about how many days can you use the sample?

OpenStudy (anonymous):

we can do this

OpenStudy (anonymous):

since the half life is 133 days you can use \[A(t)=A_0(\frac{1}{2})^{\frac{t}{133}}\]as a model for how much you have left after t days with initial amount \[A_0\] now you need 90% = 0.9 so set \[.9=(\frac{1}{2})^{\frac{t}{133}}\] solve for t via \[\frac{t}{133}=\frac{\ln(.9)}{\ln(.5)}\] \[t=\frac{133 \ln(.9)}{\ln(.5)}\] with a calculator

OpenStudy (anonymous):

whoa all wrong. after 90% is gone! leaves 10% damn sorry

OpenStudy (anonymous):

\[.1=(\frac{1}{2})^{\frac{t}{133}}\] \[\frac{t}{133}=\frac{\ln(.1)}{\ln(.5)}\] \[t=\frac{133 \ln(.1)}{\ln(.5)}=441.8\] rounded